OpenStudy (anonymous):

Identify that the two vertical lines are parallel because they both have a red arrow. This means the angles in question are alternate exterior angles, which are congruent. Your equation and solution should look like this: 8x + 14 = 9x 14 = x why is 14=x ... ? :s

OpenStudy (anonymous):

\[5-2(x-3)=2+4(x-1)-3x\]

OpenStudy (anonymous):

hi @princesshair

OpenStudy (anonymous):

8x + 14 = 9x solve for x and x=14

OpenStudy (anonymous):

8x + 14 = 9x 9x-8x = 14 x=14

OpenStudy (anonymous):

i get ittt ahhhh you guys are awesome

OpenStudy (anonymous):

That we are ;)